OpenStudy (anonymous):

Algebra Help :) Will Fan & Medal 26. Solve for x. x^2 - 4x - 5 = 0 A) x = -1 and -5 B) x = 1 and 5 C) x = 1 and -5 D) x = -1 and 5 27. Solve for x. x^2 - 10x - 13 = 0 A) 1.16, -11.16 B) -1.16, -11.16 C) 1.16, 11.16 D) 1.16, 11.16

OpenStudy (er.mohd.amir):

x*x-4x-5 =x*x-5x+x-5 =x(x-5)+1(x-5) =(x-5)(x+1)

OpenStudy (er.mohd.amir):

I is D

OpenStudy (anonymous):

the answer to 26 is D

OpenStudy (er.mohd.amir):

x*x-10x-13 use quradtatic formula x=11.16,-1.16

OpenStudy (anonymous):

and the answer to 27 is C i believe

OpenStudy (ramzy197):

26) D 27) A im not sure if you need the workings too but you can ask me if you need the workings

OpenStudy (anonymous):

@ryandaman2017 @Er.Mohd.AMIR Thanks :)